# EgoBlind: Towards Egocentric Visual Assistance for the Blind People

### v1

@@ 3.1.4 Answer Annotation
We invite 21 university students to watch the videos and answer the questions. To enhance the persuasiveness of the standard answers and reduce subjective evaluation, each question is assigned to three or four annotators for answering. Only one may see the generated answer (if available) and must verify and edit it as needed. The others are required to directly answer the question for assistance purpose. If there is no generated answer, e.g., for the manually extracted or annotated question, the annotators need to answer the question themselves. We resolve contradictory answers through consensus between the authors and the majority of annotators. Redundant answers are avoided by merging similar responses. Ultimately, each question has 1 to 4 valid answers (3 on average). Additionally, about 67.7% of the generated answers are edited by human annotators, and most unedited answers are deterministic, such as yes/no and numbers. Examples of the reference answers are shown in Tab. 1.
Figure 3: Examples of EgoBlind.
@@ 3.2 Statistic Analysis
Data Split. The annotation process takes about 10 months. We finalize 4,927 questions across 1,210 videos (two examples are shown in Fig. 3). We split the dataset into training and test sets at a 7:3 ratio while ensuring that video clips from the same source do not appear in both. As shown in Tab. 2, the training set contains 906 videos and 3,438 questions, while the test set includes 304 videos and 1,489 questions. Additionally, we employ Gemini 2.0 to probe if there is a significant discrepancy of model behaviors between the test set and the whole dataset. We find that the prediction accuracy differences are quite small, ranging from 0.3% to 0.8% across different question types. This suggests that the test set is sufficient to serve as a reliable testbed for evaluating model behaviors.
@@
906 304 1,210 3,438 1,489 4,927
(a) Distribution of question length.
(b) Distribution of answer length.
(c) Top-10 frequent questions.
(d) Top-10 frequent answers*.
@@
Figure 4: Statistic analysis of EgoBlind. *: We omit 1,123 (22.8%) “yes/no” and 541 (11.0%) “do not know” answers in (d) and (f) for better presentation. (Please zoom in for better view.)
@@
Questions. Fig. 4(a) shows that more than 80% of the questions are short in 10 words, with an average length of 7.9. The relatively shorter questions could be attributed to the use of more spoken language in practice. The word cloud in Fig. 4(e) shows that the questions are from blind users’ first-person perspective (“I”, “my”, “me”) and feature non-verbal concepts, such as “this”, “there”, “that”, “it” and “now”, which often demand on engaging in the scenarios, capturing the questioner’s hand pose and eye gaze, and reasoning the spatio-temporal contexts for understanding the questions. Fig. 4(c) shows that blind people are often interested in locating something (“where is”) and checking for potential safety issues (“are there any obstacles”).
@@
Answers. Fig. 4(b) shows that the answers are relatively longer with an average of 5.7 words. Near half of the answers have more than 3 words, with 19.4% over 10 words. A detailed study shows that the answers for tool use are significantly longer and more complex than other assistive questions; the average answer length is 12.1 for tool use while 4 to 6 for others. The most frequent answers are “yes” (13.7%), followed by “I do not know” (11.0%) and “no” (9.7%). The answer of “I do not know” is due to poor visual quality or the answer not framed by the blind users. Additionally, it may attribute to our real-time QA task setting, i.e., the answer is not visible up to the question moment. By omitting these answers, we analyze the remaining frequent answers and answer word clouds in Fig. 4(d) and Fig. 4(f), respectively. We find that the answers are mostly about directions, navigation and locations. It is worth mentioning that the locations and directions are often relative to the blind camera wearers (questioners).
@@
Videos. Fig. 4(g) shows that the vast majority of the videos (84.3%) are under 1 minute, with 15.7% exceeding 1 minute and 5.9% running longer than 2 minutes. The mean video duration is 35.8 seconds, reflecting the average duration each time a visually-impaired person recording their daily moment. Additionally, we conduct a comparison between the egocentric video content captured by sighted people (e.g., Ego4D [13] videos) and those are blind. We find that the differences are from the following major aspects: 1) Composition and Focus: Blind individuals often produce footage where subjects are off-center or out of focus (refer to Fig. 1 and Fig. 3). 2) Camera Orientation and Stability: Blind individuals hard to maintain consistent camera orientation, potentially leading to tilted or unstable footage. 3) Environmental Awareness: Blind individuals inadvertently capture obstructed views or poorly lit scenes. Nevertheless, the video resolution is high, possibly because our videos are crawled from video sharing platforms.
@@ 3.3 Dataset Comparison
The EgoBlind dataset, as a first-person perspective video question-answering dataset, has distinct uniqueness. Unlike most first-person perspective QA datasets (datasets listed in the first block of Tab. 3), which focus on assessing general-purpose visual understanding abilities, the core task of EgoBlind centers around assisting blind individuals, emphasizing the real-time and practical nature of tasks. Based on the actual needs of blind people, we categorize the questions to evaluate the model’s ability in different assistance aspects. Compared to other assist-oriented datasets (listed in the middle block of Tab. 3), EgoBlind covers the wide aspects of visual assistance for the blind people, versus episodic memory in QAEgo4D [3] or tool use in AssistQ [36] or scene-text QA in EgoTextVQA [48].
Compared to the blind-related QA datasets (such as the well-known VizWiz [14] image dataset and the recent video QA dataset VIEW-QA), EgoBlind advances VizWiz by enabling egocentric live QA with more visual information about real-world dynamics and episodic memory support. Although the videos in the VIEW-QA [33] dataset offer 360-degree visual information, the videos and questions are simulated by people with normal eyesight mimicking the perspective of blind individuals, thus lacking sufficient practical characteristics for the blind community, especially for those with long blind age. In contrast, the videos in EgoBlind are entirely filmed by blind individuals themselves, authentically reflecting their real-life scenarios. Moreover, the questions in EgoBlind are posed or verified by blind individuals of different blind ages, further ensuring the dataset’s practicality and relevance.
Table 3: Dataset Comparison. BlindV: Video captured by blind people. BlindQ: Question posed or verified by blind people. RTime: Answer are based on the video content up to the question timestamp. MA: Multiple reference answers for each question. QC: Question Category. OE/MC: Open-ended/Multi-choice.
Datasets BlindV BlindQ RTime MA QC Task
General-purpose EgoVQA
EgoVQA [10] OE
EgoTaskQA [17] OE
EgoThink [7] OE
VidEgoThink [6] OE
EgoSchema [28] MC
EgoMemoria [42] MC
Assistance EgoVQA
QAEgo4D [3] OE
AssistQ [36] MC
EgoTextVQA [48] OE
Blind VQA
VizWiz [14] - OE
VIEW-QA [33] OE
EgoBlind (Ours) OE
@@ 4 Experiment
4.1 Experimental Setups
Evaluation. We used GPT-4o mini as a validation tool to assess the model performances on EgoBlind test set. It evaluates the semantic similarity between the models’ predictions and the ground truth (reference answers). Building on previous work, we established two evaluation metrics for the models’ predictions: accuracy (0-100%, the percentage of ‘yes’ answers evaluated) and score (0-5, degree of similarity to the matching results). Specially, for each test question, a correct prediction is identified if it achieves yes-response with any one of the reference answers. It is worth mentioning that the evaluation prompts are manually finetuned to reach a maximal score agreement (0.88) between human and AI reviewers (Details are introduced in Sec. B.1). The final evaluation prompts are attached in Tab. 9.
Table 4: Performance (average accuracy and score) of sampling uniformly across the video content up to the question timestamp (video input) vs. sampling 1 frame at the question timestamp (image input). Sampling multiple frames bring better performances.
Methods Video Input Image Input
Video-LLaVA [22] 38.5 / 2.1 36.7 / 2.1
LLaMA-VID [21] 39.2 / 2.2 35.6 / 2.1
LLaVA-OV [19] 55.5 / 2.9 54.2 / 2.8
InternVL2-8B [5] 56.5 / 3.0 54.7 / 2.9
Gemini2.0[34] 56.6 / 3.0 52.4 / 2.7
Model Setup To benchmark the challenges embodied by EgoBlind, we comprehensively analyze 15 contemporary MLLMs, including 12 open-source models and 3 closed-source models (via APIs). The choice of the open-source models are based on the following criteria: 1) Achieve SOTA results on common video QA benchmarks such as ActivityNet-QA [43], NExT-QA [37], VideoChatGPT-Bench [27] and Video-MME [11]. Corresponding models are Qwen2-VL [35], LLaVA-OV [19], CogVLM2-Video [15], Video-LLaMA3 [44], InternVL-2 [5]; 2) Achieve SOTA results on recent general-purpose egocentric VideoQA benchmarks such as EgoSchema [28], VidEgoThink [6] and EgoMemoria [42]. Corresponding models are LLaVA-NeXT-Video [19], MiniCPM-V 2.6 [41] and Qwen2-VL[35]; 3) Achieve SOTA on the image blind QA dataset VizWiz [14], such as Video-LLaVA [27], LLaMA-VID[21] and VILA 1.5 [23]. We additional include ShareGPT4Video [4] as it pioneers in using Ego4D videos for pretraining.
Noteworthy, for each question, we uniformly sample the video content up to the question timestamp for predicting their answers, thus to meet the live QA task setting. Additionally, we have experimented an alternative video sampling strategy by using the frame at the question timestamp, and find that it jeopardizes the performances (refer to Tab. 4). Also, we design customized prompts (refer to Appendix Tab. 7) tailored for answering blind users’ questions for better performance (see analysis in Sec. B.2). Other details follow the official model protocols.
Table 5: Model behaviors on EgoBlind. No single model wins across all question types. Models that are superior on general-purpose egocentric VQA or image blind QA are not the best-performing. The closed-source models perform similarly to the top-performing open-source ones. All results significantly lag behind human performance.
@@ 4.1 Experimental Setups
ShareGPT4Video [4] LLaMA3-8B - 2fps 32.2 / 2.3 32.5 / 1.8 14.8 / 1.4 40.7 / 2.2 31.6 / 1.9 30.8 / 1.8 33.3 / 2.0
CogVLM2-Video [15] LLaMA3-8B 2242 24 35.6 / 2.3 42.7 / 2.3 15.7 / 1.4 50.0 / 2.6 60.1 / 3.2 33.3 / 2.0 40.0 / 2.3
Video-LLaMA3 [44] Qwen2.5-7B ori 1fps 45.6 / 2.5 54.0 / 2.7 34.8 / 2.0 48.4 / 2.6 55.3 / 2.8 56.4 / 2.9 49.6 / 2.6
InternVL2-26B [5] InternLM2-Chat-20B 4482 8 58.9 / 3.1 55.5 / 2.9 51.4 / 2.8 54.1 /2.9 47.4 / 2.6 59.8 / 3.1 54.9 / 2.9
LLaVA-OV [19] Qwen2-7B 3842 32 57.8 / 3.2 58.5 / 3.0 30.1 / 2.0 64.2 / 3.3 63.2 / 3.2 51.3 / 2.8 55.5 / 2.9
InternVL2-8B [5] InternLM2_5-7B-Chat 4482 8 61.1 / 3.3 57.2 / 3.0 48.1 / 2.7 59.3 / 3.1 47.4 / 2.5 58.1 / 3.1 56.5 / 3.0
MiniCPM-V 2.6 [41] Qwen2-7B 3842 1fps 48.9 / 2.6 47.6 / 2.5 43.3 / 2.4 32.0 / 1.9 47.4 / 2.6 45.3 / 2.4 43.0 / 2.3
Qwen2-VL [35] Qwen2-7B ori 1fps 53.3 / 3.0 50.0 / 2.6 38.6 / 2.3 35.8 / 2.0 47.4 / 2.5 50.4 / 2.7 45.1 / 2.4
LLaVA-NeXT-Video [19] Qwen2-7B 3842 1fps 41.1 / 2.4 54.8 / 2.8 35.2 / 2.1 62.0 / 3.2 52.6 / 2.9 54.7 / 2.8 52.9 / 2.8
Video-LLaVA [27] Vicuna-7B 2242 8 27.8 / 1.9 40.7 / 2.2 18.1 / 1.3 48.9 / 2.5 52.6 / 2.8 33.3 / 2.0 38.5 / 2.1
LLaMA-VID [21] Vicuna-7B 2242 1fps 35.6 / 2.2 39.8 / 2.2 30.1 / 2.0 47.8 / 2.5 50.0 / 2.7 36.8 / 2.1 39.2 / 2.2
VILA1.5 [23] LLaMA3-8B 3362 32 41.1 / 2.4 52.4 / 2.7 26.2 / 1.7 56.0 / 2.8 60.5 / 3.2 40.2 / 2.4 48.2 / 2.6
@@
Gemini 1.5 Flash - ori - 61.1 / 3.3 57.9 / 3.0 38.1 / 2.3 57.3 / 2.9 60.5 / 3.0 53.0 / 2.9 54.8 / 2.9
GPT-4o - ori 1fps 60.0 / 3.2 55.8 / 2.9 47.6 / 2.6 59.9 / 3.0 48.7 / 2.6 59.8 / 3.1 56.0 / 2.9
Gemini 2.0 - ori - 48.3 / 2.9 57.0 / 3.0 51.1 / 2.8 47.7 / 2.7 63.2 / 3.2 65.2 / 3.4 56.6 / 3.0

### v2

@@ 3.1.4 Multiple Answer Annotation
We invite 21 university students to watch the videos and answer the questions. To enhance the persuasiveness of the standard answers and reduce subjectivity, each question is assigned to three or four annotators for answering. Only one may see the generated answer (for generated questions only) and must verify and edit it as needed. The others are required to directly answer the question for assistance purpose. If there is no generated answer, e.g., for the manually extracted or annotated question, the annotators need to answer the question themselves. We resolve contradictory answers through consensus between the authors and the majority of annotators. Redundant answers are avoided by merging similar responses. Ultimately, each question has 1 to 4 valid answers (3 on average). Over 67% of the generated answers are edited by human annotators, and most unedited answers are deterministic, such as yes/no and numbers. Examples of answers are shown in Table 1.
@@ 3.2 Statistic Analysis
673 656 1,329 2,746 2,565 5,311
Data Split. The data collection and annotation process takes about 10 months. We finalize 5,311 questions across 1,329 videos. For effective and efficient evaluation, we split half of the videos (656) along with their QA pairs (2,565) as test set, while leaving the remaining for instruction tuning towards model development. The split ensures that video clips from the same source video do not appear in both sets. Additionally, we employ Gemini 2.0 to probe if there is a significant discrepancy of model behaviors between the test set and the whole dataset. We find that the prediction accuracy differences are quite small, ranging from 0.2% to 0.5% across different question types. This suggests that the test set is sufficient to serve as a reliable evaluation ground for MLLMs. The detailed numbers are presented in Table 2.
(a) Question length.
(b) Answer length.
(c) Top-10 questions.
(d) Top-10 answers*.
@@
Figure 3: Statistic analysis of EgoBlind. *: We omit 1,123 (22.6%) “yes/no” and 541 (11.0%) “do not know” answers in (d) and (f) for better presentation. (Please zoom in for better view.)
@@
Questions. Figure 3(a) shows that more than 80% of the questions are short in 10 words, with an average length of 7.9. The relatively shorter questions could be attributed to the use of more spoken language in practice. The word cloud in Figure 3(e) shows that the questions are from blind users’ first-person perspective (“I”, “me”, “my”) and feature referential concepts, such as “this”, “there”, “that”, “it” and “now”, which often demand on engaging in the scenarios, capturing the questioner’s hand motion and eye gaze, and reasoning the spatio-temporal contexts for understanding the questions. Figure 3(c) shows that blind people are often interested in locating something (“where is”) and checking for potential safety issues (“are there any obstacles”).
@@
Answers. Figure 3(b) shows that the answers are relatively longer with an average of 5.7 words. Near half of the answers have more than 3 words, with 19.4% over 10 words. A detailed study shows that the answers for tool use are significantly longer and more complex than other assistive questions; the average answer length is 12.1 for tool use while 4 to 6 for others. The most frequent answers are “yes” (13.0%), followed by “I do not know” (10.7%) and “no” (9.6%). The answer of “I do not know” is due to poor visual quality or the answer not framed by the blind users. Additionally, it may be attributed to our real-time QA task setting, i.e., the answer is not visible up to the question moment. We keep these unanswerable questions to evaluate if models can reject to answer rather than hallucinate potentially malicious answers. For better analysis, we omit these answers and analyze the remaining frequent answers and answer word clouds in Figure 3(d) and 3(f), respectively. We find that the answers are mostly about directions, navigation and locations. It is worth mentioning that the locations and directions are often relative to the camera wearers (questioners).
@@
Videos. Figure 3(g) shows that the vast majority of the videos (82.3%) are within 1 minute, with 17.7% exceeding 1 minute and 7.1% running longer than 2 minutes. The mean video duration is 40.0 seconds, reflecting the average duration each time a visually-impaired person recording their daily moment. Additionally, we conduct a comparison between the egocentric video content captured by sighted people (e.g., Ego4D [32] videos) and those are blind (examples are presented in Appendix A.3). We find that the differences are from the following major aspects: 1) Composition and Focus: Blind individuals often produce footage where subjects are off-center or out of focus (refer to Figure 1 and Appendix). 2) Camera Orientation and Stability: Blind individuals hard to maintain consistent camera orientation, potentially leading to tilted or unstable footage. 3) Environmental Awareness: Blind individuals inadvertently capture obstructed views or poorly lit scenes. Nevertheless, the video resolution is higher, possibly because the videos are crawled from video sharing platforms.
@@ 3.3 Dataset Comparison
Unlike most Ego-VQA datasets (the first block of Table 3), which focus on general-purpose visual understanding, EgoBlind emphasizes assisting blind individuals in different assistance aspects. Compared to other assist-oriented datasets (the middle block), EgoBlind covers the wide aspects of visual assistance for the blind people, versus episodic memory in QAEgo4D [35] or tool use in AssistQ [31] or scene-text QA in EgoTextVQA [38].
Compared to the blind-related QA datasets (the bottom block), EgoBlind advances VizWiz [24] by enabling egocentric live QA with more visual information about real-world dynamics and episodic memory support. Although the videos in the VIEW-QA [13] dataset offer 360-degree visual information, the videos and questions are simulated by singted people mimicking blind individuals, thus lacking sufficient practical characteristics for the blind community, especially for those with long blind age. In contrast, the videos in EgoBlind are entirely filmed by blind individuals themselves, authentically reflecting their real-life scenarios. Moreover, the questions in EgoBlind are posed or verified by blind individuals of different blind ages, further ensuring the dataset’s practicality and relevance. Additionally, EgoBlind supports a live QA setting with timestamp annotation of each question and the answers are based on the video contents prior to the question timestamp.
Table 3: Dataset Comparison. BV: Video captured by blind people. BQ: Question posed or verified by blind people. RT: Answe based on the video content up to the question timestamp. MA: Multiple reference answers for each question. QC: Question Category. OE/MC: Open-ended/Multi-choice.
Datasets # V # Q VLen(s) BV BQ RT MA QC Task
General Purpose EgoTaskQA [9] 2K 40K 25.0 OE
VidEgoThink [10] 217 600 23.4 OE
EgoSchema [11] 5K 5K 180.0 MC
EgoMemoria [20] 629 7K 858.5 MC
Assist QAEgo4D [35] 166 1.8K 495.1 OE
AssistQ [31] 100 531 115.0 MC
EgoTextVQA(In) [38] 813 2,2K 101.7 OE
EgoTextVQA(Out) [38] 694 4.8K 10.0 OE
Blind VizWiz [24] - - - OE
VIEW-QA [13] 1.0K 4.1K 34.4 OE
EgoBlind (Ours) 1.3K 5.3K 40.0 OE
@@ 4 Experiment
4.1 Experimental Setup
Evaluation. We used GPT-4o mini as a validation tool to assess the model performances on EgoBlind test set. It evaluates the semantic similarity between the models’ predictions and the ground truth (reference answers). Building on previous work, we established two evaluation metrics for the models’ predictions: accuracy (0-100%, the percentage of ‘yes’ answers evaluated) and score (0-5, degree of similarity to the matching results). Specially, for each question, a correct prediction is identified if it achieves yes-response with any one of the reference answers. Note that the evaluation prompts are manually finetuned to reach a maximal score agreement (0.88) between human and AI reviewers (details in Appendix B.1).
Model Setup To benchmark the challenges carried by EgoBlind, we comprehensively analyze 15 contemporary MLLMs, including 12 open-source models and 3 closed-source models (via APIs). The choice of the open-source models are based on the following criteria: 1) Achieve SOTA results on common video QA benchmarks such as NExT-QA [8], VideoChatGPT-Bench [18] and Video-MME [19]. Corresponding models are LLaVA-OV [15], CogVLM2-Video [47], Video-LLaMA3 [16], InternVL2.5 [14]; 2) Achieve SOTA results on general-purpose egocentric VideoQA benchmarks such as EgoSchema [11], VidEgoThink [10] and EgoMemoria [20]. Corresponding models are LLaVA-Video [48], MiniCPM-V 2.6 [49] and Qwen2.5-VL[4]; 3) Achieve SOTA on the image blind QA dataset VizWiz [24], such as Video-LLaVA [18], LLaMA-VID[22] and VILA 1.5 [23]. Noteworthy, for each question, we uniformly sample the video content up to the question timestamp for answer prediction, thus to match the live QA task setting. Moreover, we design customized prompts tailored for answering blind users’ questions for better performance. Additionally, we obtain human performance by inviting 3 university students who did not participant in annotation.
Table 4: Model behaviors (accuracy / score) on EgoBlind.
@@ 4.1 Experimental Setup
ShareGPT4Video [46] LLaMA3-8B - 2fps 36.0 / 2.3 34.5 / 2.0 18.6 / 1.5 43.5 / 2.3 40.3 / 2.2 26.0 / 1.7 33.9 / 2.0
CogVLM2-Video [47] LLaMA3-8B 2242 24 37.4 / 2.4 45.8 / 2.4 15.5 / 1.4 51.4 / 2.6 44.4 / 2.4 34.1 / 2.0 41.2 / 2.3
Video-LLaMA3 [16] Qwen2.5-7B ori 1fps 46.9 / 2.5 54.5 / 2.8 38.9 / 2.2 53.6 / 2.8 48.6 / 2.6 53.8 / 2.8 51.3 / 2.7
LLaVA-OV [15] Qwen2-7B 3842 16 59.9 / 3.1 57.1 / 3.0 31.4 / 2.1 65.9 / 3.3 58.3 / 3.0 50.2 / 2.7 55.1 / 2.9
InternVL2.5-8B [14] InternLM2_5-Chat-7B 4482 8 59.2 / 3.2 57.9 / 3.0 43.8 / 2.6 60.1 / 3.0 44.4 / 2.4 54.3 / 2.9 55.8 / 2.9
InternVL2.5-26B [14] InternLM2_5-Chat-20B 4482 8 69.4 / 3.4 58.6 / 3.0 49.5 / 2.7 57.0 / 2.9 43.1 / 2.3 53.2 / 2.8 56.6 / 2.9
MiniCPM-V 2.6 [49] Qwen2-7B 3842 1fps 51.0 / 2.5 49.8 / 2.6 38.4 / 2.2 34.2 / 1.9 37.5 / 2.0 42.8 / 2.2 43.6 / 2.3
Qwen2.5-VL [4] Qwen2.5-7B ori 1fps 56.5 / 3.1 53.8 / 2.8 31.2 / 2.1 53.5/ 2.8 41.7 / 2.4 38.2 / 2.2 49.0 / 2.7
LLaVA-Video [48] Qwen2-7B 3842 1fps 39.5 / 2.2 54.7 / 2.8 34.8 / 2.1 61.9 / 3.1 48.6 / 2.6 49.7 / 2.6 52.0 / 2.7
Video-LLaVA [18] Vicuna-7B 2242 8 21.8 / 1.8 42.4 / 2.3 21.4 / 1.5 48.5 / 2.5 41.7 / 2.2 36.4 / 2.0 39.1 / 2.2
LLaMA-VID [22] Vicuna-7B 2242 1fps 33.3 / 2.3 41.8 / 2.3 19.6 / 1.5 52.2 / 2.7 40.3 / 2.2 37.0 / 2.1 40.1 / 2.2
VILA1.5 [23] LLaMA3-8B 3362 8 45.6 / 2.6 49.6 / 2.6 25.3 / 1.6 58.5 / 2.9 50.0 / 2.6 41.0 / 2.3 47.2 / 2.5
@@
Gemini 1.5 Flash - ori - 74.2 / 3.7 59.4 / 3.0 44.3 / 2.6 60.4 / 3.0 58.3 / 2.9 54.3 / 2.9 57.9 / 3.0
Gemini 2.0 Flash - ori - 57.8 / 3.0 58.8 / 3.0 48.4 / 2.7 56.1 / 2.8 52.8 / 2.7 48.6 / 2.6 55.7 / 2.9
Gemini 2.5 Flash - ori - 66.7 / 3.1 57.9 / 3.0 50.3 / 2.7 55.1 / 2.8 48.6 / 2.5 58.4 / 3.0 56.3 / 2.9
GPT-4o - ori 1fps 65.3 / 3.4 64.7 / 3.3 51.0 / 2.9 58.8 / 3.0 62.4 / 3.1 55.5 / 2.9 60.6 / 3.1